OpenStudy (logan13):

What is the value of the expression? (5 + 10) ÷ 5

OpenStudy (goose.):

Parrentheses are first, so 5+10= 15 then you divide 15 by 5 which gives you 3.

OpenStudy (imtiaz7):

First you need to solve the brackets 5+10=15 and then 15 divided by 5 that is : 15/5=3